# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 10.1 Application to Bernoulli Randomized Design
::=:= $\mathbb{E}\left[{\vec{X}}=(\Delta_{g},\Theta_{gl},\ldots,\Theta_{g1},\mathcal{T}_{g},\Lambda_{g},\Delta_{h},\Theta_{hl},\ldots,\Theta_{h1},\mathcal{T}_{h},\Lambda_{h})^{\top}\right].$
@@
where $\vec{{\cal Y}}(N)$ denotes the sample mean of observed outcomes over time and across various subpopulations, corresponding to the vector on the left-hand side of (95). The vector $\vec{{\cal B}}(N)$ represents the unknown coefficients on the right-hand side of (95), while ${\cal X}(N)$ denotes a $b(T-l+1)$ by $21+l+2+l+$ 2 matrix. This matrix is aligned with the vectors on the right-hand side of (95), but with the corresponding sample means of observed outcomes and treatments replacing the asymptotic terms, see (97).
@@
Suppose that ${\cal X}(N)^{\top}{\cal X}(N)$ is invertible. Let ^$\vec{\widehat{{\cal B}}}(N):=\big{(}{\cal X}(N)^{\top}{\cal X}(N)\big{)}^{-1}{\cal X}(N)^{\top}\vec{{\cal Y}}(N)$ be the least squares estimator. Then, ^$\vec{\widehat{{\cal B}}}(N)$ provides a strongly consistent estimator for the coefficient vector $(\delta,\theta_{gl},\ldots,\theta_{g1},\tau_{g},\lambda_{g},\theta_{hl},\ldots,\theta_{h1},\tau_{h},\lambda_{h})^{\top}$.
Proof. Because the matrix ${\cal X}(N)^{\top}{\cal X}(N)$ is invertible, the estimator ^$\vec{\widehat{{\cal B}}}(N)$ is a continuous function of the input data. Therefore, we can pass the limit through the estimator function as:
@@ 10.2 First-order Estimators
Given subpopulation $\mathcal{S}$, Algorithms 4 and 5 aim to estimate counterfactuals under the desired treatment allocation over $\mathcal{S}$ using $b$ distinct subpopulations $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$ as the estimation samples. Both algorithms share their first two steps. In the first step, they compute sample means of observed outcomes for both the entire population and each subpopulation, along with sample means of delivered and desired treatment allocations. The second step estimates unknown parameters in the state evolution equation (93) using least squares estimation as detailed in Proposition 10.5. The algorithms then diverge in their third step, applying these results through two distinct approaches detailed below. The consistency proofs for both algorithms follow directly from earlier results and are omitted for brevity.
@@ 10.3 Higher-order Recursive Estimators
Estimate $\bm{\Theta}_{g}\in\mathbb{R}^{m\times n_{1}}$ and $\bm{\Theta}_{h}\in\mathbb{R}^{m\times n_{2}}$ as $\widehat{\bm{\Theta}}_{g}$ and $\widehat{\bm{\Theta}}_{h}$:
@@ 11 Detrending for Temporal Patterns
While semi-recursive estimators (see §10.2.1) can capture complex temporal patterns in unit outcomes, they face two major limitations. First, they cannot estimate out-of-sample counterfactuals because their architecture relies directly on observed sample means. Second, unlike recursive estimators (see Algorithm 6), they cannot accommodate higher-order approximations of the outcome functions ($g_{t}$ and $h_{t}{}$). This limitation stems from their dependence on the closed-form of state evolution equation (as outlined in (93)), which may not exist for more complex characterizations of the outcome functions.
@@ 11.2 Augmented Causal Message-Passing Model
Additionally, $\tilde{g}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\tilde{h}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\vec{0}$, and functions $\tilde{g}_{t}$ and $\tilde{h}_{t}$ satisfy the conditions of Assumption 9.4. Note that enforcing conditions $\tilde{g}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\tilde{h}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\vec{0}$ ensures that $\vec{Y}_{t+1}{}(\mathbf{0})=\vec{y}^{\;}_{t+1}{}$, aligning with (98). Letting $\vec{{\cal Y}}_{t}{}(\bm{W}{}):=\vec{Y}_{t}{}(\bm{W}{})-\vec{y}^{\;}_{t}{}$, we can then rewrite the augmented model (99) to match the dynamics of the original outcome specification:
@@
We emphasize that Equations (98) and (100) provide distinct characterizations of the outcomes, and Algorithm 7 requires both to hold simultaneously. Specifically, assuming the conditions of §9.4 hold in both settings, the baseline outcomes $\vec{y}^{\;}_{t}{}$ satisfy the state evolution equation corresponding to (98). However, in the context of (100), state evolution becomes relevant only when treatment is delivered (i.e., $\bm{W}{}\neq\mathbf{0}$). Indeed, under Assumption 11.2, the third condition in Assumption 9.4 indicates that state evolution can be derived only when there exists a non-zero treatment effect.
@@ 12 Auxiliary Results
Let $\left\{X_{n,i}:1\leq i\leq n,\;n\geq 1\right\}$ be a triangular array of random variables such that $(X_{n,1},\ldots,X_{n,n})$ are mutually independent with a mean equal to zero for each $n$ and $\frac{1}{n}\sum_{i=1}^{n}E\left[|X_{n,i}|^{2+\kappa}\right]\leq cn^{\kappa/2}$ for some $10<\kappa<$ 1 and $c<\infty$. Then, we have
@@
We also need the following form of the law of large numbers which is an extension of Lemma 4 in Bayati and Montanari, (2011).
@@
[…]

### v2

@@ 11.1 Application to Bernoulli Randomized Design
= $\mathbb{E}\left[{\vec{X}}=(\Delta_{g},\Theta_{gl},\ldots,\Theta_{g1},\mathcal{T}_{g},\Lambda_{g},\Delta_{h},\Theta_{hl},\ldots,\Theta_{h1},\mathcal{T}_{h},\Lambda_{h})^{\top}\right].$
@@
where $\vec{{\cal Y}}(N)$ denotes the sample mean of observed outcomes over time and across various subpopulations, corresponding to the vector on the left-hand side of (102). The vector $\vec{{\cal B}}(N)$ represents the unknown coefficients on the right-hand side of (102), while ${\cal X}(N)$ denotes a b(Tl+1)b(T-l+1) by $1+l+2+l+2$ matrix. This matrix is aligned with the vectors on the right-hand side of (102), but with the corresponding sample means of observed outcomes and treatments replacing the asymptotic terms, see (104).
@@
Suppose that ${\cal X}(N)^{\top}{\cal X}(N)$ is invertible. Let $\vec{\widehat{{\cal B}}}(N):=\big({\cal X}(N)^{\top}{\cal X}(N)\big)^{-1}{\cal X}(N)^{\top}\vec{{\cal Y}}(N)$ be the least squares estimator. Then, $\vec{\widehat{{\cal B}}}(N)$ provides a strongly consistent estimator for the coefficient vector $(\delta,\theta_{gl},\ldots,\theta_{g1},\tau_{g},\lambda_{g},\theta_{hl},\ldots,\theta_{h1},\tau_{h},\lambda_{h})^{\top}$.
Proof. Because the matrix ${\cal X}(N)^{\top}{\cal X}(N)$ is invertible, the estimator $\vec{\widehat{{\cal B}}}(N)$ is a continuous function of the input data. Therefore, we can pass the limit through the estimator function as:
@@ 11.2 First-order Estimators
Given subpopulation $\mathcal{S}$, Algorithms 5 and 6 aim to estimate counterfactuals under the desired treatment allocation over $\mathcal{S}$ using $b$ distinct subpopulations $\mathcal{S}_{1},\ldots,\mathcal{S}_{b}$ as the estimation samples. Both algorithms share their first two steps. In the first step, they compute sample means of observed outcomes for both the entire population and each subpopulation, along with sample means of delivered and desired treatment allocations. The second step estimates unknown parameters in the state evolution equation (100) using least squares estimation as detailed in Proposition 11.5. The algorithms then diverge in their third step, applying these results through two distinct approaches detailed below. The consistency proofs for both algorithms follow directly from earlier results and are omitted for brevity.
@@ 11.3 Higher-order Recursive Estimators
Estimate $\bm{\Theta}_{g}\in\mathbb{R}^{m\times n_{1}}$ and $\bm{\Theta}_{h}\in\mathbb{R}^{m\times n_{2}}$, respectively, as $\widehat{\bm{\Theta}}_{g}$ and $\widehat{\bm{\Theta}}_{h}$:
@@ 12 Detrending for Temporal Patterns
While semi-recursive estimators (see §11.2.1) can capture complex temporal patterns in unit outcomes, they face two major limitations. First, they cannot estimate out-of-sample counterfactuals because their architecture relies directly on observed sample means. Second, unlike recursive estimators (see Algorithm 7), they cannot accommodate higher-order approximations of the outcome functions ($g_{t}$ and $h_{t}$). This limitation stems from their dependence on the explicit form of the state evolution equation (as outlined in (100)), which may not exist for more complex characterizations of the outcome functions.
@@ 12.2 Augmented Causal Message-Passing Model
Additionally, $\tilde{g}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\tilde{h}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\vec{0}$, and functions $\tilde{g}_{t}$ and $\tilde{h}_{t}$ satisfy the conditions of Assumption 8. Note that enforcing conditions $\tilde{g}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\tilde{h}_{t}\left(\vec{0},\mathbf{0},\bm{X}\right)=\vec{0}$ ensures that $\vec{Y}_{t+1}{}(\mathbf{0})=\vec{y}^{\;}_{t+1}{}$, aligning with (105). Letting $\vec{{\cal Y}}_{t}{}(\bm{W}{}):=\vec{Y}_{t}{}(\bm{W}{})-\vec{y}^{\;}_{t}{}$, we can then rewrite the augmented model (106) to match the dynamics of the original outcome specification:
@@
We emphasize that Equations (105) and (107) provide distinct characterizations of the outcomes, and Algorithm 8 requires both to hold simultaneously. Specifically, assuming the conditions of §10.4 hold in both settings, the baseline outcomes $\vec{y}^{\;}_{t}{}$ satisfy the state evolution equation corresponding to (105). However, in the context of (107), state evolution becomes relevant only when treatment is delivered (i.e., $\bm{W}{}\neq\mathbf{0}$). Indeed, under Assumption 12.2, the third condition in Assumption 8 indicates that state evolution can be derived only when there exists a non-zero treatment effect.
@@ 13 Auxiliary Results
Let $\left\{X_{n,i}:1\leq i\leq n,\;n\geq 1\right\}$ be a triangular array of random variables such that $(X_{n,1},\ldots,X_{n,n})$ are mutually independent with a mean equal to zero for each $n$ and $\frac{1}{n}\sum_{i=1}^{n}E\left[|X_{n,i}|^{2+\kappa}\right]\leq cn^{\kappa/2}$ for some $0<\kappa<1$ and $c<\infty$. Then, we have
@@
We also need the following form of the law of large numbers, which is an extension of Lemma 4 in Bayati and Montanari, (2011), and the proof can be found in the work of Shirani and Bayati, (2024).
@@
$W_{1}\left(\mathcal{L}\left(\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{n\in\mathcal{S}}\Phi^{n}\right),{\cal N}\left(0,\frac{1}{N}\right)\right)=\sup\left\{\mathbb{E}\left[f\left(\frac{1}{\sqrt{\left|{\mathcal{S}}\right|}}\sum_{n\in\mathcal{S}}\Phi^{n}\right)\right]-\mathbb{E}\left[f\left(\frac{Z}{\sqrt{N}}\right)\right]\;\Bigg|\;f\text{ is 1-Lipschitz}\right\},$ (110)